OpenStudy (anonymous):

Simplify the following expression Square root sign -144 times square root sign -48

OpenStudy (anonymous):

\[\sqrt{-144}\sqrt{-48}\]

OpenStudy (anonymous):

\[\sqrt{-144} \times \sqrt{-48}\]

OpenStudy (anonymous):

\[\sqrt{144i^2}\times \sqrt{48i^2}\]

OpenStudy (anonymous):

\[12i \times4\sqrt{3}i\]

OpenStudy (anonymous):

\[-48\sqrt{3}\]

OpenStudy (anonymous):

thank you that's what i got too i just wasn't sure about it

OpenStudy (anonymous):

yr welcome :)